Project

General

Profile

03262014 Minutes

Attending: Michael, Evan, Mat, Gavin, Pavan

Operation:
Beam off for two days
DAQ testing - Large system instability, and RC
APD installs on diblock continue
Diblock 8 cooling delayed until next week

P1-
NUMI and cosmic triggers are down, and number of hits/spill down from 3/25 18:00 to 20:00 yesterday. (DDT)
P2-
P2 ran out of sync for a few hours evidenced in slice metrics. Used the incorrect configuration.
P3-
Humidity in Ash River should be fixed according to Bill.

NearDet - Patern mode is still running.
APD - Are several weeks out?
FEB soon - suggesting to add FEB stuff into P3

Watchdog plot - Adding a vertical at 25 run.

Nearline reprocessing? - Wait for the final release.

Work toward DQ blessing plots.